1 Experience indicated that the firm to build 20 devices would use 400 hours to build the first device. According to the past experience, the time will be shortened with the increase of the number of machines to be built. In the bid, the company is assumed the learning rate to be 90% and the skilled labour cost to be $80 per hour. Find the followings. e) The company wants to have the time taken to produce the 20th units to be exactly 220 hours. What the learning rate (L) should be to meet the company’s wish?

GIVE THE ANSWER IN DETAIL
Answer #1

Answer e :

Given thta,

The company wants to have the time taken to produce, the 20th units to be exactly 220 hours :

Y20 = Y1 * 20^{LN(r) / LN(2)} = 220

LN(r) / LN(2) = (LN(220) - LN(400))/LN(20)

LN(r) = LN(2) * (LN(220) - LN(400))/LN(20) = -0.1383

r = EXP(-0.1383) = 87%

The learning rate (L) should be to meet the company’s wish = 87%